# Generic ETH: Eigenstate Thermalization beyond the Microcanonical

### v1

@@ 1 Introduction
In the case of a closed quantum system, the Eigenstate Thermalization Hypothesis (ETH)Srednicki:1994mfb ; Deutsch:1991msp ; Jensen:1984gu provides the prevailing framework for addressing the question of which operators and states will thermalize in a given system. Most directly, the ETH is an ansatz for the form of operator matrix elements in the energy eigenbasis, along with the suggestion that it is such operators that thermalize, at least in states that have narrow energy support. The ETH has been well-supported both numerically and analytically in various settings rigol2008thermalization ; mondaini2016eigenstate , and applied to domains ranging from conformal field theory Lashkari:2016vgj ; Dymarsky:2019etq to holography Pollack:2020gfa ; Saad:2019pqd ; Altland:2021rqn . Nevertheless, open questions still remain regarding its domain of applicability. An informal, but well-motivated expectation is that it is the local and otherwise “simple” operators in chaotic systems which take the ETH form. By contrast, integrable systems (here meaning systems with an extensive number of conserved charges) generally do not thermalize and correspondingly have simple observables which are not well-characterized by ETH. In fact, this is not a strict dichotomy. Many states in integrable systems have been shown to relax to a Generalized Gibbs Ensemble Rigol_Dunjko_Yurovsky_PhysRevLett.98.050405 ; Cramer_Dawson_Eisert_Osborne_PhysRevLett.100.030602 which take account of conserved parameters additional to the energy. In such cases, a generalized eigenstate thermalization hypothesis cassidy_2011 ; He_Santos_Wright_Rigol_2013 ; Rigol:2016itf offers an explanation for this behavior again in terms of operator matrix elements. Once again, the necessary and sufficient conditions for the occurrence of generalized ETH, and the extent to which it suffices to account for all instances of thermalization or lack thereof, is the subject of ongoing research. Various additional generalizations of standard ETH have also been considered, for instance to the case of noncommuting charges Murthy:2022dao , and in another direction to account for higher moment correlations in operator coefficients Foini_Kurchan_2019 . In this work, we do not attempt a complete review of these topics, but we do intend to provide a gentle introduction for the uninitiated, replete with examples in lattice systems.
@@ 2.2 Eigenstate Thermalization Hypothesis
where $\bar{E}$ is the average energy $\frac{1}{2}(E_{i}+E_{j})$ and $\omega:=E_{i}-E_{j}$ the energy difference. This form is expected to hold for many local and physically relevant operators in chaotic systems, at least out to near the edges of the spectrum. Importantly, all off-diagonal elements are exponentially suppressed, with $S(\bar{E})$ representing the thermal entropy at energy $\bar{E}$. This can be thought of as the density of states at energy $\bar{E}$, or the number of active energy eigenstates states in a thermal (Gibbs) state at energy expectation $\bar{E}$, scaling proportionally to the system size. More generally, the entropic suppression could be written as $S(\bar{E},\omega)$, but it is common to restrict $\omega$-dependence to the “envelope function” $f_{\mathcal{O}}(\bar{E},\omega)$. This function is largely unspecified and context dependent, except that it is expected to fall off with increasing $|\omega|$ and becomes flat for $|\omega|<\hbar/\tau^{*}$, where $\tau^{*}$ is the longest physical timescale associated with the system, often set by the (inverse) Thouless energy JTEdwards_1972 in diffusive systems.
@@ 2.3 Connection to random matrix theory and quantum chaos
To further understand the connection between random matrix theory and thermalization, we here briefly summarize a few basic results from this field. This also allows us to motivate the definition of quantum chaos employed in this work. See Forrester:2003um for a succinct historical overview of applications of random matrix theory in physics, along with relevant citations. See Kota2014book , for instance, for an introduction to the properties of common ensembles, which we will now only briefly introduce.
@@
Aside from the empirical connection noted in the Berry-Tabor and BGS conjectures, the use of level-spacing statistics to define quantum chaos seems to bear little resemblance to classical definitions, which are usually based on the rapid divergence of nearby phase space trajectories. Parallels can be drawn, however. Level-repulsion is an avoidance of degeneracies, which can be thought of as an avoidance of conserved (or approximately conserved) quantities. This is commensurate with the classical dichotomy between chaotic and integrable systems. The same level-repulsion helps ensure that chaotic systems undergo rapid and uniform dephasing, which is crucial for thermalization. Despite these qualitative connections, the occurence of a Wigner-Dyson distribution of nearest-neighbor eigenvalue spacings is far from a complete characterization of a system’s spectrum, and it occurs in systems with much more structure than the Gaussian random ensembles. These ensembles describe a class of Hamiltonians that are, in some sense, “as chaotic as possible”, but the extent to which the onset of chaos in terms of level-spacings alone guarantees other behaviors traditionally associated with chaotic systems, and appearance of the ETH form of local operators, is a question that we will probe in the remainder of this work.
@@ 2.4 Generic ETH
Let us briefly discuss the effect of the fluctuations of diagonal terms on the ETH equilibrium prediction, captured in the second term in equation (11). If the $c_{i}$ and $R_{ii}$ are uncorrelated, the sum vanishes on average, while its variance is described by $\overline{\sum_{i}|c_{i}|^{4}|R_{ii}|^{2}|f(E_{i},0)|^{2}e^{-S(E_{i})}}$. The overbar denotes an average over instantiations of $c_{i}$, thus to define it rigorously would require specifying some ensemble of states to consider. Regardless, the entropic suppression serves to keep the sum small. Likewise, if the $c_{i}$ are nonzero across some number of states $N_{\Psi}$, with an average squared magnitude of $1/N_{\Psi}$, then increasing $N_{\Psi}$ can also serve to reduce these fluctuations as $1/N_{\Psi}$. This happens in any fixed microcanonical window as the density of states increases with system size, but also by broadening the support outside a microcanonical window at fixed system size. If the support moves closer to the edge of the spectrum (where the density of states drops), the entropic suppression is less effective, but the effect of including more states may still be suppressive.
@@ 3.4 Thermalization in more detail
For the system considered here with the standard/chaotic coefficients, the Hamiltonian itself is the only quasi-local conserved charge, so we expect the energy (or temperature) to be the only parameter in the relevant thermal ensemble. In order to test thermalization across a range of energies/temperatures, we employ random unentangled states with variable energy expectation values. By randomly preparing unentangled states (see footnote 12), the energy of the resulting states tends to be close to zero, the center of the spectrum. We can, however, choose to keep only those states that fall near some other target energy $E:=\bra{\Psi}\hat{H}\ket{\Psi}$. A corresponding Gibbs ensemble
@@
Figure 4: Illustrations of the improving thermalization with increasing system size $L$. In the left panel, the blue curve is a thermal expectation value for the first qubit entanglement entropy in a system of size $L=12$, which we take as representative of larger systems. In both plots, red datapoints of increasing opacity represent data for increasing system size, from $L=4$ to $L=10$, interpolated for clarity. In the left (right) panel, each point corresponds to the average (standard deviation) of the first qubit entanglement entropy, sampled over a large late time window ($\Delta t\sim 1000$). The resulting average (standard deviation) was then again averaged over 300 initially-unentangled states at the corresponding energy expectation value.
@@ 3.5 ETH in the qubit spin chain
We test these expectations in figure 5. The left panel 4(a) focuses on the first point. In every eigenstate, the expectation value of the local operator $\sigma_{x}^{(1)}$ as well as the entropy on the first qubit are plotted in comparison to the thermal predictions at the same energy (from the Gibbs state with corresponding energy expectation value). The conformity between the eigenstates and thermal predictions is apparent, though it becomes weaker towards the edges of the spectrum.
@@
The right panels 4(b) both serve to illustrate the suppression of off-diagonal matrix elements with increasing system size. The top panel considers the ratio of the average diagonal element magnitude to that of off-diagonal elements for two operators: $\text{Ratio}:=\overline{|\text{diag}|}\big{/}\overline{|\text{off-diag}|}$. For the local operator, $\sigma_{x}^{(1)}$, this ratio is seen to increase exponentially with system size. By contrast the operator $\prod_{i}^{L}\sigma_{x}^{(i)}$, which has the same spectrum but is supported along the entire spin chain, shows a ratio approximately independent of system size. To compare more explicitly with the rate of exponential suppression appearing as $\exp{-S(\bar{E})/2}$ ) in the ETH ansatz, we do a second test in the bottom panel of the figure. Instead of averaging over all off-diagonals, we average specifically over the counter-diagonal, meaning we compute
@@
It is interesting to note that the random unentangled states are not the paradigmatic states of narrow energy support often used in studies of ETH and thermalization. States with narrow energy support are particularly natural to consider in quench scenarios, where one starts from a system eigenstate and then moves out of equilibrium through a sudden modification of Hamiltonian parameters. But ETH is predictive for a much broader set of states. The fact that thermalization occurs for the states considered here might at first seem surprising, but the reason in simple. Having randomly generated unentangled states and then selected those falling at a particular energy, the resulting energy distribution tends to mimic that of a Gibbs distribution, which is qualitatively illustrated in figure 7. This figure compares energy distributions from a selection of random unentangled states to energy density distributions of a corresponding Gibbs state. [^15] The smooth Gibbs state energy distributions in the right hand side of figure 7 are obtained by assuming a Wigner semicircle density of states, which is applicable to GOE ensemble Hamiltonians, scaled to the same energy range as the qubit Hamiltonian. The actual density of states for the qubit Hamiltonian in fact has more features than the ideal Wigner semicircle, but we did not create any smooth approximation to it. Thus the comparison between left and right figures is merely qualitative. It is clear that the energy distribution of these states is far from a microcanonical state, and much closer to a Gibbs state.
@@ 5 Conclusions
Throughout this work, we have emphasized that the (generalized) ETH form of simple operators, where it occurs, is more predictive than just its most common application to (generalized) microcanonical states. For a very large class of states, it clarifies that deviations between local observables’ equilibrium behavior and the predictions of thermal distributions are best captured simply by modifying the latter to mimic higher central moments of the states’ energy and charge distributions. For observables with approximately linear dependence on microcanonical thermal parameters, thermalization occurs with exceptional insensitivity to the details of the initial state, which can be broadly supported well outside microcanonical windows. We dub this behavior ‘generic ETH,’ and illustrate with a few simple, but natural sets of states in our lattice systems.
With our focus on exact diagonalization and direct demonstration of eigenstate thermalization in these systems, many interesting phenomenological aspects of the designed qutrit system have been left unexplored. It would be interesting to see in more detail how the time scales involved in thermalization are affected by the presence of the quasilocal charge. The manner in which a chaotic qubit system is embedded into an enlarged lattice with an overlaid local charge dynamics may lead to interesting interplay of multiple timescales, associated with the qubit system dynamics and the charge spreading separately. Since the size of the state space that is “active” on the chaotic qubit subsystem depends directly on the amount of charge on a lattice site, there may be a rich phenomenology not just of thermalization and entanglement growth, but also other measures of information scrambling and chaos not studied in this work. It would also be interesting to allow for local “frustrations” of the charge spreading ability that partially segregate otherwise uniformly chaotic Hamiltonian dynamics. All these possibilities we leave to future work.

### v2

@@ 1 Introduction
In the case of a closed quantum system, the Eigenstate Thermalization Hypothesis (ETH)Srednicki:1994mfb ; Deutsch:1991msp ; Jensen:1984gu provides the prevailing framework for addressing the question of which operators and states will thermalize in a given system. Most directly, the ETH is an ansatz for the form of operator matrix elements in the energy eigenbasis, along with the suggestion that it is such operators that thermalize, at least in states that have narrow energy support. The ETH has been well-supported both numerically and analytically in various settings rigol2008thermalization ; mondaini2016eigenstate , and applied to domains ranging from conformal field theory Lashkari:2016vgj ; Dymarsky:2019etq to holography Pollack:2020gfa ; Saad:2019pqd ; Altland:2021rqn . Nevertheless, open questions still remain regarding its domain of applicability. An informal, but well-motivated expectation is that it is the local and otherwise “simple” operators in chaotic systems which take the ETH form. By contrast, integrable systems (here meaning systems with an extensive number of conserved charges) generally do not thermalize and correspondingly have simple observables which are not well-characterized by ETH. In fact, this is not a strict dichotomy. Many states in integrable systems have been shown to relax to a Generalized Gibbs Ensemble Rigol_Dunjko_Yurovsky_PhysRevLett.98.050405 ; Cramer_Dawson_Eisert_Osborne_PhysRevLett.100.030602 which take account of conserved parameters additional to the energy. In such cases, a generalized eigenstate thermalization hypothesis cassidy_2011 ; He_Santos_Wright_Rigol_2013 ; Rigol:2016itf offers an explanation for this behavior again in terms of operator matrix elements. Once again, the necessary and sufficient conditions for the occurrence of generalized ETH, and the extent to which it suffices to account for all instances of thermalization or lack thereof, is the subject of ongoing research in varied contexts Dymarsky:2019etq ; PhysRevLett.131.060401 ; Pappalardi:2023nsj . Various additional generalizations of standard ETH have also been considered, for instance to the case of noncommuting charges Murthy:2022dao , and in another direction to account for higher moment correlations in operator coefficients Foini_Kurchan_2019 . In this work, we do not attempt a complete review of these topics, but we do intend to provide a gentle introduction for the uninitiated, replete with examples in lattice systems.
@@ 2.2 Eigenstate Thermalization Hypothesis
where $\bar{E}$ is the average energy $\frac{1}{2}(E_{i}+E_{j})$ and $\omega:=E_{i}-E_{j}$ the energy difference. This form is expected to hold for many local and physically relevant operators in chaotic systems, at least out to near the edges of the spectrum. Importantly, all off-diagonal elements are exponentially suppressed, with $S(\bar{E})$ representing the thermal entropy at energy $\bar{E}$. This can be thought of as the density of states at energy $\bar{E}$, or the number of active energy eigenstates states in a thermal (Gibbs) state at energy expectation $\bar{E}$, scaling proportionally to the system size. More generally, the entropic suppression could be written as $S(\bar{E},\omega)$, but it is common to restrict $\omega$-dependence to the “envelope function” $f_{\mathcal{O}}(\bar{E},\omega)$. This function is largely unspecified and context dependent. It is expected to fall off with increasing $|\omega|$ and becomes flat for $|\omega|<\hbar/\tau^{*}$, where $\tau^{*}$ is a system timescale beyond which random matrix theory results apply [^4] In systems with conserved charges, this is often a time associated with diffusion across the full system, called the Thouless time JTEdwards_1972 . See Friedman:2019gyi for analysis of this timescale in a Floquet system with conserved charge, and Roy:2020tgl ; Roy:2022zig ; Kumar:2023jvb for analysis in periodically driven many body systems with and without a $U(1)$ symmetry for fermionic, bosonic, and mixed species, respectively. . The detailed form of $f_{\mathcal{O}}(\bar{E},\omega)$, particularly its falloff with $|\omega|$, has been related to the decay of nonequal-time correlation functions and linear response relaxation times after perturbations PhysRevLett.111.050403 . In Sorokhaibam:2022tgq it is found that consistency with the second law of thermodynamics constraints the generic form of $f_{\mathcal{O}}(\bar{E},\omega)$ to monotonically increase with $S(\bar{E})$, in the sense that this leads to nondecreasing entropy when $\mathcal{O}$ is a perturbing operator.
@@ 2.3 Connection to random matrix theory and quantum chaos
To further understand the connection between random matrix theory and thermalization, we here briefly summarize a few basic results from this field. This also allows us to motivate the definition of quantum chaos employed in this work. See Forrester:2003um for a succinct historical overview of applications of random matrix theory in physics, along with relevant citations. See Kota2014book , for instance, for an introduction to the properties of common matrix ensembles, which we will now only briefly introduce.
@@
Aside from the empirical connection noted in the Berry-Tabor and BGS conjectures, the use of level-spacing statistics to define quantum chaos seems to bear little resemblance to classical definitions, which are usually based on the rapid divergence of nearby phase space trajectories. Qualitative parallels can be drawn, however. Level-repulsion is an avoidance of degeneracies, which can be thought of as an avoidance of conserved (or approximately conserved) quantities. This is commensurate with the classical dichotomy between chaotic and integrable systems. The same level-repulsion helps ensure that chaotic systems undergo rapid and uniform dephasing, as is crucial for thermalization. More rigorously, recent work Vikram:2022xiq has connected spectral statistics to the dynamical criteria of “quantum cyclic ergodicity and aperiodicity” (notions which generalize classical ergodicity on a discretized phase space to arbitrary quantum systems), finding that systems with Wigner-Dyson statistics satisfy both criteria. Of course, the occurence of a Wigner-Dyson distributions in nearest-neighbor eigenvalue spacings is far from a complete characterization of a system’s spectrum; it occurs in systems with much more structure than the Gaussian random ensembles, which represent Hamiltonians that are in some sense “as chaotic as possible”. Further characterizations of the spectrum can be found in quantities such as the spectral form factor Mehta1967book ; Brezin:1997rze , which is essentially the Fourier transform of the two point correlation function of the eigenvalue distribution, and has been studied in a wide variety of systems (for example, see Cotler:2016fpe ; Gharibyan:2018jrp ; Friedman:2019gyi ). The extent to which any given characterisation of chaos, such as the onset of Wigner-Dyson level-spacings employed in this work, guarantees other behaviors traditionally associated with chaotic systems, and appearance of the ETH form of simple operators, is a question that we will probe in the remainder of this work.
@@ 2.4 Generic ETH
Let us briefly discuss the effect of the fluctuations of diagonal terms of the ETH ansatz. When the ETH fails, and nonvanishing fluctuations remain in thermodynamic limit for diagonal elements of local operators, this has been directly linked to the presence of (quasi)local integrals of motion in the system PhysRevLett.124.040603 . On the other hand, if the ETH ansatz applies, the affect of these terms on the equilibrium prediction in a finite system is captured in the second term in equation (11). If the $c_{i}$ and $R_{ii}$ are uncorrelated, the sum vanishes on average, while its variance is described by $\overline{\sum_{i}|c_{i}|^{4}|R_{ii}|^{2}|f(E_{i},0)|^{2}e^{-S(E_{i})}}$. The overbar denotes an average over instantiations of $c_{i}$, thus to define it rigorously would require specifying some ensemble of states to consider. Of course, the entropic suppression serves to keep the sum small. Likewise, if the $c_{i}$ are nonzero across some number of states $N_{\Psi}$, with an average squared magnitude of $1/N_{\Psi}$, then increasing $N_{\Psi}$ can also serve to reduce these fluctuations as $1/N_{\Psi}$. This happens in any fixed microcanonical window as the density of states increases with system size, but also by broadening the support outside a microcanonical window at fixed system size. If the support moves closer to the edge of the spectrum (where the density of states drops), the entropic suppression is less effective, but the effect of including more states may still be suppressive. This will be evident in the thermalization of broadly-supported (non-microcanonical) states shown in later examples, e.g. in figure 16.
@@ 3.4 Thermalization in more detail
For the system considered here with the standard/chaotic coefficients, the Hamiltonian itself is the only quasi-local conserved charge, so we expect the energy (or temperature) to be the only parameter in the relevant thermal ensemble. In order to test thermalization across a range of energies/temperatures, we employ random unentangled states with variable energy expectation values. By randomly preparing unentangled states (see footnote 13), the energy of the resulting states tends to be close to zero, the center of the spectrum. We can, however, choose to keep only those states that fall near some other target energy $E:=\bra{\Psi}\hat{H}\ket{\Psi}$. A corresponding Gibbs ensemble
@@
Figure 4: Illustrations of the improving thermalization with increasing system size $L$. In the left panel, the blue curve is a thermal expectation value for the first qubit entanglement entropy in a system of size $L=12$, which we take as representative of larger systems. In both plots, red datapoints of increasing opacity represent data for increasing system size, from $L=4$ to $L=10$, interpolated for clarity. In the left (right) panel, each point corresponds to the average (standard deviation) of the first qubit entanglement entropy, sampled over a large late time window ($\Delta t\sim 1000$). The resulting average (standard deviation) was then averaged over 300 initially-unentangled states at the corresponding energy expectation value.
@@ 3.5 ETH in the qubit spin chain
We test these expectations in figure 5. The left panel 5(a) focuses on the first point. In every eigenstate, the expectation value of the local operator $\sigma_{x}^{(1)}$ as well as the entropy on the first qubit are plotted in comparison to the thermal predictions at the same energy (from the Gibbs state with corresponding energy expectation value). The conformity between the eigenstates and thermal predictions is apparent, though it becomes weaker towards the edges of the spectrum.
@@
The right panels 5(b) both serve to illustrate the suppression of off-diagonal matrix elements with increasing system size. The top panel considers the ratio of the average diagonal element magnitude to that of off-diagonal elements for two operators: $\text{Ratio}:=\overline{|\text{diag}|}\big{/}\overline{|\text{off-diag}|}$. For the local operator, $\sigma_{x}^{(1)}$, this ratio is seen to increase exponentially with system size. By contrast the operator $\prod_{i}^{L}\sigma_{x}^{(i)}$, which has the same spectrum but is supported along the entire spin chain, shows a ratio approximately independent of system size. To compare more explicitly with the rate of exponential suppression appearing as $\exp{-S(\bar{E})/2}$ ) in the ETH ansatz, we do a second test in the bottom panel of the figure. Instead of averaging over all off-diagonals, we average specifically over the counter-diagonal, meaning we compute
@@
It is interesting to note that the random unentangled states are not the paradigmatic states of narrow energy support often used in studies of ETH and thermalization. States with narrow energy support are particularly natural to consider in quench scenarios, where one starts from a system eigenstate and then moves out of equilibrium through a sudden modification of Hamiltonian parameters. But ETH is predictive for a much broader set of states. The fact that thermalization occurs for the states considered here might at first seem surprising, but the reason in simple. Having randomly generated unentangled states and then selected those falling at a particular energy, the resulting energy distribution tends to mimic that of a Gibbs distribution, which is qualitatively illustrated in figure 7. This figure compares energy distributions from a selection of random unentangled states to energy density distributions of a corresponding Gibbs state. [^16] The smooth Gibbs state energy distributions in the right hand side of figure 7 are obtained by assuming a Wigner semicircle density of states, which is applicable to GOE ensemble Hamiltonians, scaled to the same energy range as the qubit Hamiltonian. The actual density of states for the qubit Hamiltonian in fact has more features than the ideal Wigner semicircle, thus the comparison between left and right figures is merely qualitative. It is clear that the energy distribution of these states is far from a microcanonical state, and much closer to a Gibbs state.
@@ 5 Conclusions
Throughout this work, we have emphasized that the (generalized) ETH form of simple operators, where it occurs, is more predictive than just its most common application to (generalized) microcanonical states. For a large class of states with non-negligible fluctuations in conserved quantities, it clarifies that deviations between local observables’ equilibrium behavior and the predictions of thermal distributions are best captured simply by modifying the latter to mimic higher central moments of the states’ energy and charge distributions. For observables with approximately linear dependence on microcanonical thermal parameters, thermalization occurs with exceptional insensitivity to the details of the initial state, which can be broadly supported well outside microcanonical windows. We dub this behavior ‘generic ETH,’ and illustrate with a few simple, but natural sets of states in our lattice systems.
With our focus on exact diagonalization and direct demonstration of eigenstate thermalization in these systems, many interesting phenomenological aspects of the designed qutrit system have been left unexplored. It would be interesting to see in more detail how the time scales involved in thermalization are affected by the presence of the quasilocal charge. The manner in which a chaotic qubit system is embedded into an enlarged lattice with an overlaid local charge dynamics may lead to interesting interplay of multiple timescales, associated with the chaotic qubit system dynamics and the charge diffusion separately. Since the size of the state space that is “active” on the chaotic qubit subsystem depends directly on the amount of charge on a lattice site, there may be a rich phenomenology not just of thermalization and entanglement growth, but also other measures of information scrambling and chaos not studied in this work. It would also be interesting to allow for local “frustrations” of the charge spreading ability that partially segregate otherwise homogeneous chaotic dynamics. All these possibilities we leave to future work.